How much energy is required to heat 87.1 g acetone (molar mass=58.08 g/mol) from a solid at -154.0°C to a liquid at -42.0°C? The following physical data may be useful.
1
Expert's answer
2020-02-14T04:36:49-0500

Moles of acetone in 87.1 grams of acetone="\dfrac{given\ mass }{molar\ mass}=\dfrac{87.1}{58}=1.5 moles"


Std enthalpy change of fusion of acetone , "\u0394H^o_{fus}=5.7 kJ\/mol"

Melting point of acetone is -95 0C

Energy change in transition from -1540 C to -95 0C = "nC_p\Delta T=1.5*96*59=8496 J"

Energy change in phase change i.e. solid to liquid="n*\Delta H^0_{fus}=1.5*5.7=8.55 kJ=8550J"

Energy change in transition from -950C to -420C = "nC_p\Delta T=1.5*125.5*53=9977 J"


"Total\ energy\ required = Sum \ of\ all\ the \ above\ energy = 8496+8550+9977 = 27023 J=27.023 kJ"